# A Data-Driven Aggressive Autonomous Racing Framework Utilizing Local Trajectory Planning with Velocity Prediction

### v1

@@ Abstract
The development of autonomous driving has boosted the research on autonomous racing. However, existing local trajectory planning methods have difficulty planning trajectories with optimal velocity profiles at racetracks with sharp corners, thus weakening the performance of autonomous racing. To address this problem, we propose a local trajectory planning method that integrates Velocity Prediction based on Model Predictive Contour Control (VPMPCC). The optimal parameters of VPMPCC are learned through Bayesian Optimization (BO) based on a proposed novel Objective Function adapted to Racing (OFR). Specifically, VPMPCC achieves velocity prediction by encoding the racetrack as a reference velocity profile and incorporating it into the optimization problem. This method optimizes the velocity profile of local trajectories, especially at corners with significant curvature. The proposed OFR balances racing performance with vehicle safety, ensuring safe and efficient BO training. In the simulation, the number of training iterations for OFR-based BO is reduced by 42.86% compared to the state-of-the-art method. The optimal simulation-trained parameters are then applied to a real-world F1TENTH vehicle without retraining. During prolonged racing on a custom-built racetrack featuring significant sharp corners, the mean velocity of VPMPCC reaches 93.18% of the vehicle’s handling limits. The released code is available at https://github.com/zhouhengli/VPMPCC.
@@ I INTRODUCTION
Autonomous racing stems from the development of autonomous driving. Pushing vehicle performance to its limits is the goal of autonomous racing. This poses a significant challenge to local trajectory planning methods due to the following four factors: (a) minimum lap time guarantee [1, 2, 3] ; (b) control feasibility of the planned trajectory[4, 5]; (c) computational efficiency of the optimization problem[6, 7]; and (d) safe and efficient parameters tuning[8, 9, 10]. Global trajectory planning[1, 11] leverages complete racetrack information for racing trajectory planning. However, it fails to account for real-time vehicle state changes, causing control issues and thus reducing racing performance. Recently, local trajectory planning methods integrating planning and control have been widely studied[12, 13, 14, 15]. Model Predictive Contour Control (MPCC) is a typical example of such an approach[13, 14]. To realize minimum lap time, MPCC introduces the projected velocity as an effective metric to represent the racing progress. This method performs well on racetracks without significant curvature corners[14]. However, with sharp corners on the racetrack (as shown in Fig. 1), relying solely on tracking error to balance the projected velocity maximization makes it difficult for MPCC to plan cornering trajectories with optimal velocities[16]. To tackle this problem, we propose a Velocity Prediction MPCC (VPMPCC) local trajectory planning method. Specifically, the proposed VPMPCC realizes velocity prediction by using vehicle longitudinal velocity as an independent velocity decision variable. The key advantage lies in optimizing the velocity profile of the local trajectory, particularly on corners with significant curvature, ultimately reducing lap time.
@@
In recent years, data-driven Bayesian optimization (BO)[8, 17, 9] has been developed to realize automatic parameters tuning. However, these methods do not consider modeling vehicle safety while reducing lap time, resulting in the inability to guarantee vehicle safety during generalization[18]. Meanwhile, simultaneously ensuring the accuracy of the vehicle model and achieving computational efficiency of the optimization problem is also a primary challenge[5, 6]. To bridge these gaps, a safe and efficient data-driven racing framework is proposed, incorporating the VPMPCC and BO. Specifically, in closed-loop BO training, the proposed framework uses the high-fidelity vehicle dynamics model obtained from offline system identification to filter the racing trajectory planned by the vehicle kinematic model-based VPMPCC. Moreover, a practical Objective Function adapted to Racing (OFR) is designed for BO to achieve efficient training. The advantage of this framework is that the parameters of the VPMPCC planner can be safely trained in the simulation and transferred to real-world vehicles without retraining. Therefore, the vehicle performance can be pushed to its limits efficiently and safely. In summary, this paper provides the following contributions:
@@
    A data-driven autonomous racing framework safely pushes the vehicle to its handling limits, validated through well-established and reliable real-world vehicle experiments. The VPMPCC and BO-based autonomous racing framework (VPBO-RF) demonstrates superior performance on a self-built 1:10 scale F1TENTH vehicle, excelling on a challenging racetrack with several sharp corners. Despite limited computational resources and execution delays, the mean velocity of VPMPCC achieved 93.18% of vehicle handling limits, with an efficient mean computation time of 7.04 ms.
This paper consists of several sections. Section II introduces MPCC and the system dynamic of VPMPCC, followed by an overview of BO. Section III introduces the implementation of VPMPCC. Section IV describes the overall framework of VPBO-RF, focusing on the design of the OFR. Section V includes simulations and real-world experiments, while Section VI serves as the conclusion, summarizing the essential findings and outlining directions for future research.
@@ II Preliminaries
This section includes an introduction to MPCC and the system dynamics of VPMPCC, followed by an overview of conventional BO.
II-A Model Predictive Contouring Control
Figure 2: Schematic diagram illustrating the calculation of contour and lag errors in conventional MPCC. The proposed VPMPCC method incorporates the vehicle’s longitudinal velocity as an independent decision variable.
@@ II-A Model Predictive Contouring Control
As shown in Fig. 2, the contour error $e_{\textrm{con}}$ and the lag error $e_{\textrm{lag}}$ of MPCC[13] are calculated as follows:
@@
$f_{\textrm{MPCC}}=-\sum_{k=1}^{N_{p}}\frac{\gamma\cdot v_{{\textrm{p}},k}}{v_{\textrm{max}}}+\|\mathbf{c}\cdot\mathbf{e}_{s_{k}}\|^{2}_{\mathbf{q}_{\mathbf{e}}}+\lVert\mathbf{u}_{k+1}-\mathbf{u}_{k}\rVert^{2}_{\mathbf{R}}$ (2)
where $N_{p}$ denotes the prediction horizon and $\gamma$ is the scalar weight of maximum projected velocity. The weight matrix $\mathbf{q}_{\mathbf{e}}$ for contour and lag errors is $\mathrm{diag}\left(\begin{bmatrix}q_{\mathbf{e}_{\textrm{con}}}&q_{\mathbf{e}_{\textrm{lag}}}\end{bmatrix}\right)$, and the weight matrix $\mathbf{R}$ is for the change of control variables. $v_{\textrm{max}}$ is the maximum longitudinal velocity used for normalization.
II-B System Dynamics
In contrast to the conventional MPCC, VPMPCC decouples the projected velocity $v_{\textrm{p}}$ from the vehicle longitudinal velocity $v$. The advantage of this design is that the introduction of velocity prediction can be a trade-off with maximizing the projected velocity at corners with significant curvature. Therefore, the system dynamics used by VPMPCC is:
where $\varphi$ is the yaw angle and $L$ is the wheelbase. Then let $\bm{\zeta}=[x,\,y,\,\varphi,\,s]^{\top}$ denotes the state space, and $\mathbf{u}=[v,\,\delta,\,v_{\textrm{p}}]^{\top}$ denotes the configuration space. The final control inputs for the underlying chassis are $v$ and steering angle $\delta$. Therefore, the weight matrix $\mathbf{R}$ in (2) is expressed as $\mathrm{diag}\left(\begin{bmatrix}q_{\Delta v}&q_{\Delta\delta}&{q}_{\Delta v_{\textrm{p}}}\end{bmatrix}\right)$.
@@ III Velocity Prediction MPCC
This section outlines the implementation of VPMPCC, emphasizing velocity prediction, and discusses the trade-off between velocity prediction and maximizing projected velocity.
@@ III-A Design of the Velocity Prediction
The VPMPCC is proposed to optimize the velocity profile for local trajectory planning on racetracks with sharp curvature corners. The velocity prediction is accomplished by enhancing the similarity between planned velocities in the prediction horizon and the corresponding reference velocity in RVP, which is parameterized by the arc length as ${v_{\textrm{RVP},{s}}}$. Therefore, the following objective function for velocity prediction can be constructed:
$f_{\textrm{VP}}$ $=\frac{q_{v}}{v_{\textrm{max}}}\sum_{k=1}^{N_{p}}\left(v_{k}-{v_{\textrm{RVP},{s_{k}}}}\right)^{2}$ (5)
where $v_{\Delta,\textrm{max}}$ denotes the maximum error in velocity matching. The $q_{v}$ is a scalar weight. VPMPCC uses the minimum curvature path of [11] as the reference line and the corresponding velocity profile as the RVP. The VPMPCC controls the racing performance through the velocity prediction component $f_{\textrm{VP}}$ and the maximizing projected velocity component in the (2) together. At racetrack corners with high curvature, maximizing the projected velocity component incurs a significant cost in the velocity prediction component. This forces the optimization problem to shift from the aggressive decision of maximizing the vehicle’s projected velocity to the safer decision of reducing the velocity. It is important to note that the RVP is used only for velocity prediction, not velocity tracking. This distinction is the fundamental difference between VPMPCC and longitudinal velocity tracking controllers[20].
@@ III-B VPMPCC Formation
where $\bm{\zeta}_{\textrm{max}}$ and $\bm{\zeta}_{\textrm{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\textrm{max}}$ and $\mathbf{u}_{\textrm{min}}$ are the upper and lower limits of the control variable, respectively. $T_{s}$ is the prediction interval, and $\bm{\zeta}_{\textrm{cur}}$ is current state values. The racetrack width is scaled using $\kappa<1$, ensuring the planned trajectory remains fully within the racetrack for safety.
@@ IV Data-driven Autonomous Racing Framework Utilizing High-efficiency OFR-driven BO
This section outlines the OFR design, focusing on racing performance and vehicle safety modeling, followed by the method for filtering abnormal trajectories using the vehicle dynamics model. The overall training framework is shown in Fig. 3. The vehicle dynamics model used during BO training and the system identification method are detailed in [21]. The parameters of the proposed VPMPCC to be optimized are expressed in the following form:
$\bm{\theta}={\begin{bmatrix}N_{p}&q_{v}&\gamma&e_{\textrm{con}}&e_{\textrm{lag}}&q_{\Delta v}&q_{\Delta\delta}&{q}_{\Delta v_{\textrm{p}}}&\kappa\end{bmatrix}}^{\top}_{9\times 1}$
where the meaning of the parameters in $\bm{\theta}$ is described in detail in Section II and  III. The observations chosen for BO is $\mathbf{z}=\{T_{\textrm{lap}},\,\mathbf{d},\,\mathbf{P}\}$. $T_{\textrm{lap}}$ denotes lap time. $\mathbf{d}=[d_{1},\,\cdots,\,d_{N_{\mathbf{z}}}]$ is the orthogonal distance of the racing trajectory corresponding to the reference line. $\mathbf{P}=[\mathbf{p}_{1},\,\cdots,\,\mathbf{p}_{N_{\mathbf{z}}}]$ is the list of the racing trajectory points. $N_{\mathbf{z}}$ is the number of trajectory points. The number of training episodes is $N_{\textrm{BO}}$.
@@ IV-1 Surrogate model
The OFR denoted as $J_{\bm{\theta}}$ consists of the following components: lap time, trajectory length minimization, and limitations on orthogonal distance. Firstly, the modeling of ${L(\mathbf{z})}$ encourages VPMPCC to explore lower lap time is:
@@
where $\lambda_{2},\lambda_{3}$ achieves a trade-off between different components. $D_{\textrm{ref}}$ is the length of the reference line. The purpose of this component is to force VPMPCC to search for shorter racing trajectories, thereby accelerating the exploration of minimal lap time.
@@
The vehicle kinematics model cannot fully capture the dynamics of the vehicle, leading to abnormal trajectories during racing, as shown in Fig. 3. This is because control inputs calculated based on the vehicle kinematics model may lead to abnormal position shifts after being integrated over time by the vehicle dynamics model. During BO training, these abnormal trajectories are identified as unqualified and penalized through ${J}_{\textrm{fail}}$, preventing BO from further exploring the unqualified trajectories generated by the kinematic model. Therefore, the $J_{\bm{\theta}}$ can be expressed as:
@@
where $\Delta\mathbf{p}=\|\mathbf{p}_{i}-\mathbf{p}_{i+1}\|,\,\forall i\in[1,...,N_{\mathbf{z}-1}]$. The $d_{\textrm{ub}}$ denotes the maximum distance between neighboring points on the racing trajectory, and $D_{\textrm{lb}}$ denotes the minimum racing trajectory length. The value of OFR for crashed and collision racing trajectories are all ${J}_{\textrm{fail}}$.
@@ IV-2 Acquisition function
The acquisition function chosen is expected improvement[22], which is represented as:
@@
where $[a]^{+}=\max{(a,0)}$. And $\tilde{J}_{n}^{*}$ is best value to date. The next evaluation point $\bm{\theta}_{i+1}$ is then selected using $\bm{\theta}_{i+1}=\arg\max_{\bm{\theta}\in\Theta}\alpha_{\textrm{EI}}(\bm{\theta};\mathcal{D}_{n})$. VPMPCC is then reconstructed based on $\bm{\theta}_{i+1}$ for collecting new observations $\mathbf{z}_{i+1}$, which is used to enrich $\mathcal{D}_{n}$, thus completing the closed-loop training.

### v2

@@ Abstract
The development of autonomous driving has boosted the research on autonomous racing. However, existing local trajectory planning methods have difficulty planning trajectories with optimal velocity profiles at racetracks with sharp corners, thus weakening the performance of autonomous racing. To address this problem, we propose a local trajectory planning method that integrates Velocity Prediction based on Model Predictive Contouring Control (VPMPCC). The optimal parameters of VPMPCC are learned through Bayesian Optimization (BO) based on a proposed novel Objective Function adapted to Racing (OFR). Specifically, VPMPCC achieves velocity prediction by encoding the racetrack as a reference velocity profile and incorporating it into the optimization problem. This method optimizes the velocity profile of local trajectories, especially at corners with significant curvature. The proposed OFR balances racing performance with vehicle safety, ensuring safe and efficient BO training. In the simulation, the number of training iterations for OFR-based BO is reduced by 42.86% compared to the state-of-the-art method. The optimal simulation-trained parameters are then applied to a real-world F1TENTH vehicle without retraining. During prolonged racing on a custom-built racetrack featuring significant sharp corners, the mean projected velocity of VPMPCC reaches 93.18% of the vehicle’s handling limits. The released code is available at https://github.com/zhouhengli/VPMPCC.
@@ I INTRODUCTION
Autonomous racing stems from the development of autonomous driving. Pushing vehicle performance to its limits is the goal of autonomous racing. This poses a significant challenge to local trajectory planning methods due to the following four factors: (a) minimum lap time guarantee [1, 2, 3] ; (b) control feasibility of the planned trajectory[4, 5, 6]; (c) computational efficiency of the optimization problem[7, 8, 9]; and (d) safe and efficient parameters tuning[10, 11, 12]. Global trajectory planning[1, 13] leverages complete racetrack information for racing trajectory planning. However, it fails to account for real-time vehicle state changes, causing control issues and thus reducing racing performance. Recently, local trajectory planning methods integrating planning and control have been widely studied[14, 15, 16, 17]. Model Predictive Contouring Control (MPCC) is a typical example of such an approach[15, 16]. To realize minimum lap time, MPCC introduces the projected velocity as an effective metric to represent the racing progress. This method performs well on racetracks without significant curvature corners[16]. However, with sharp corners on the racetrack (as shown in Fig. 1), relying solely on tracking error to balance the projected velocity maximization makes it difficult for MPCC to plan cornering trajectories with optimal velocities[18]. To tackle this problem, we propose a Velocity Prediction MPCC (VPMPCC) local trajectory planning method. Specifically, the proposed VPMPCC realizes velocity prediction by using vehicle longitudinal velocity as an independent velocity decision variable. The key advantage lies in optimizing the velocity profile of the local trajectory, particularly on corners with significant curvature, ultimately reducing lap time.
@@
In recent years, data-driven Bayesian optimization (BO)[10, 11] has been developed to realize automatic parameters tuning. However, these methods do not consider modeling vehicle safety while reducing lap time, resulting in the inability to guarantee vehicle safety during generalization[19]. Meanwhile, simultaneously ensuring the accuracy of the vehicle model and achieving computational efficiency of the optimization problem is also a primary challenge[5, 7]. To bridge these gaps, a safe and efficient data-driven racing framework is proposed, incorporating the VPMPCC and BO. Specifically, in closed-loop BO training, the proposed framework uses the high-fidelity vehicle dynamics model obtained from offline system identification to filter the racing trajectory planned by the vehicle kinematic model-based VPMPCC. Moreover, a practical Objective Function adapted to Racing (OFR) is designed for BO to achieve efficient training. The advantage of this framework is that the parameters of the VPMPCC planner can be safely trained in the simulation and transferred to real-world vehicles without retraining. Therefore, the vehicle performance can be pushed to its limits efficiently and safely. In summary, this paper provides the following contributions:
@@
    A data-driven autonomous racing framework safely pushes the vehicle to its limits, validated through reliable real-world experiments. The VPMPCC and BO-based autonomous racing framework (VPBO-RF) outperform on a self-built 1:10 scale F1TENTH vehicle, excelling on a challenging racetrack with sharp corners. Despite limited computational resources and execution delays, the mean projected velocity of VPMPCC achieved 93.18% of vehicle handling limits, with an efficient mean computation time of 7.04 ms.
This paper consists of several sections. Section II introduces the system dynamic of VPMPCC and conventional MPCC, followed by an overview of BO. Section III introduces the implementation of VPMPCC. Section IV describes the overall framework of VPBO-RF, focusing on the design of the OFR. Section V includes simulations and real-world experiments, while Section VI serves as the conclusion, summarizing the essential findings and outlining directions for future research.
@@ II Preliminaries
This section presents the system dynamics of VPMPCC, an introduction to MPCC, and an overview of BO.
II-A System Dynamics
In contrast to the conventional MPCC, VPMPCC decouples the projected velocity $v_{\textrm{p}}$ from the vehicle longitudinal velocity $v$. The advantage of this design lies in the introduction of velocity prediction, enabling a trade-off between maximizing the projected velocity, thus handling corners with significant curvature. Therefore, the system dynamics used by VPMPCC is:
where $\varphi$ is the yaw angle and $L$ is the wheelbase. Then let $\bm{\zeta}=[x,\,y,\,\varphi,\,s]^{\top}$ denote the state space, and $\mathbf{u}=[v,\,\delta,\,v_{\textrm{p}}]^{\top}$ denotes the configuration space. The final control inputs for the underlying chassis are $v$ and steering angle $\delta$. Therefore, the weight matrix $\mathbf{R}$ in (3) is expressed as $\mathrm{diag}\left(\begin{bmatrix}q_{\Delta v}&q_{\Delta\delta}&{q}_{\Delta v_{\textrm{p}}}\end{bmatrix}\right)$.
II-B Model Predictive Contouring Control
Figure 2: Schematic diagram illustrating the calculation of contouring and lag errors in conventional MPCC. The proposed VPMPCC method incorporates the vehicle’s longitudinal velocity as an independent decision variable.
@@ II-B Model Predictive Contouring Control
As shown in Fig. 2, the contouring error $e_{\textrm{con}}$ and the lag error $e_{\textrm{lag}}$ of MPCC[15, 14] are calculated as follows:
@@
$f_{\textrm{MPCC}}=\sum_{k=1}^{N_{p}}-\frac{\gamma\cdot v_{{\textrm{p}},k}}{v_{\textrm{max}}}\cdot T_{s}+\|\mathbf{c}\cdot\mathbf{e}_{s_{k}}\|^{2}_{\mathbf{q}_{\mathbf{e}}}+\lVert\mathbf{u}_{k+1}-\mathbf{u}_{k}\rVert^{2}_{\mathbf{R}}$ (3)
where $N_{p}$ denotes the prediction horizon and $\gamma$ is the scalar weight of maximum projected velocity. The weight matrix $\mathbf{q}_{\mathbf{e}}$ for contouring and lag errors is $\mathrm{diag}\left(\begin{bmatrix}q_{\mathbf{e}_{\textrm{con}}}&q_{\mathbf{e}_{\textrm{lag}}}\end{bmatrix}\right)$, and the weight matrix $\mathbf{R}$ is for the change of control variables. $v_{\textrm{max}}$ is the maximum longitudinal velocity used for normalization.
@@ III Velocity Prediction MPCC
This section outlines the VPMPCC implementation, emphasizing velocity prediction and the trade-off between velocity prediction and maximizing projected velocity.
@@ III-A Design of the Velocity Prediction
The VPMPCC is proposed to optimize the velocity profile for local trajectory planning on racetracks with sharp curvature corners. The velocity prediction is accomplished by enhancing the similarity between planned velocities in the prediction horizon and the corresponding reference velocity in RVP. Therefore, the following cost function for velocity prediction can be constructed:
$f_{\textrm{VP}}$ $=\frac{q_{v}}{v_{\Delta,\textrm{max}}}\sum_{k=1}^{N_{p}}\left(v_{k}-{v_{\textrm{RVP},{s_{k}}}}\right)^{2}$ (5)
where $v_{\Delta,\textrm{max}}$ denotes the maximum error in velocity matching. The ${v_{\textrm{RVP},{s_{k}}}}$ represents the parameterized RVP based on arc length, and $k$ denotes the time instance within prediction horizon. The $q_{v}$ is a scalar weight. VPMPCC uses the minimum curvature path of [13] as the reference line and the corresponding velocity profile as the RVP. The VPMPCC controls the racing performance through the velocity prediction component $f_{\textrm{VP}}$ and the maximizing projected velocity component in the Eq. (3) together. At racetrack corners with high curvature, maximizing the projected velocity component incurs a significant cost in the velocity prediction component. This forces the optimization problem to shift from the aggressive decision of maximizing the vehicle’s projected velocity to the safer decision of reducing the velocity. It is important to note that the RVP is used only for velocity prediction, not velocity tracking. This distinction is the fundamental difference between VPMPCC and longitudinal velocity tracking controllers[21].
@@ III-B VPMPCC Formation
where $\bm{\zeta}_{\textrm{max}}$ and $\bm{\zeta}_{\textrm{min}}$ denote the upper and lower bounds of the state constraints, while $\mathbf{u}_{\textrm{max}}$ and $\mathbf{u}_{\textrm{min}}$ are the upper and lower limits of the control variable, respectively. $T_{s}$ is the prediction interval, and $\bm{\zeta}_{\textrm{cur}}$ is current state values. The racetrack width is scaled using $\alpha<1$, ensuring the planned trajectory remains fully within the racetrack for safety.
@@ IV Data-driven Autonomous Racing Framework Utilizing High-efficiency OFR-driven BO
This section outlines the OFR design, focusing on racing performance and vehicle safety modeling, followed by the method for filtering abnormal trajectories using the vehicle dynamics model. The overall training framework is shown in Fig. 3. The vehicle dynamics model and the system identification method are detailed in [22]. The parameters of the proposed VPMPCC to be optimized are:
$\bm{\theta}={\begin{bmatrix}N_{p}&q_{v}&\gamma&e_{\textrm{con}}&e_{\textrm{lag}}&q_{\Delta v}&q_{\Delta\delta}&{q}_{\Delta v_{\textrm{p}}}&\alpha\end{bmatrix}}^{\top}_{9\times 1}$
where the meaning of the parameters in $\bm{\theta}$ is described in detail in Section II and  III. The observations chosen for BO is $\mathbf{z}=\{T_{\textrm{lap}},\,\mathbf{d},\,\mathbf{P}\}$. $T_{\textrm{lap}}$ denotes lap time. $\mathbf{d}=[d_{1},\,\cdots,\,d_{N_{\mathbf{z}}}]$ is the orthogonal distance of the racing trajectory from the reference line. $\mathbf{P}=[\mathbf{p}_{1},\,\cdots,\,\mathbf{p}_{N_{\mathbf{z}}}]$ is the list of the racing trajectory points. $N_{\mathbf{z}}$ is the number of trajectory points. The number of training episodes is $N_{\textrm{BO}}$.
@@ IV-1 Surrogate model
The OFR denoted as $J_{\bm{\theta}}$ consists of the following components: lap time, trajectory length minimization, and limitations on orthogonal distance. Firstly, the modeling of ${L(\mathbf{z})}$ encouraging VPMPCC to explore lower lap time is:
@@
where $\lambda_{2},\lambda_{3}$ achieve a trade-off between different components. $D_{\textrm{ref}}$ is the length of the reference line. The purpose of this component is to force VPMPCC to search for shorter racing trajectories, thereby accelerating the exploration of minimal lap time.
@@
The vehicle kinematics model cannot fully capture the dynamics of the vehicle, leading to abnormal trajectories during racing, as shown in Fig. 3. This is because control inputs calculated based on the vehicle kinematics model may lead to abnormal position shifts after being integrated over time by the vehicle dynamics model. During BO training, these abnormal trajectories are identified as unqualified and penalized through ${J}_{\textrm{fail}}$, preventing further exploration of these trajectories. Therefore, the $J_{\bm{\theta}}$ can be expressed as:
@@
where $\Delta\mathbf{p}=\|\mathbf{p}_{i}-\mathbf{p}_{i+1}\|,\,\forall i\in[1,...,N_{\mathbf{z}-1}]$. The $d_{\textrm{ub}}$ denotes the maximum distance between neighboring points on the racing trajectory, and $D_{\textrm{lb}}$ denotes the minimum racing trajectory length. These metrics are designed to penalize trajectories with position drifts. The value of OFR for crashed racing trajectories is all ${J}_{\textrm{fail}}$.
@@ IV-2 Acquisition function
The acquisition function chosen is Expected Improvement (EI), which is represented as:
@@
where $[a]^{+}=\max{(a,0)}$, and $\tilde{J}_{n}^{*}$ is best value to date. The Matèrn kernel with parameter $\nu=\frac{5}{2}$ is used as the kernel function[12]. The next evaluation point $\bm{\theta}_{i+1}$ is then selected using $\bm{\theta}_{i+1}=\arg\max_{\bm{\theta}\in\Theta}\alpha_{\textrm{EI}}(\bm{\theta};\mathcal{D}_{n})$. VPMPCC is then reconstructed based on $\bm{\theta}_{i+1}$ for collecting new observations $\mathbf{z}_{i+1}$, which is used to enrich $\mathcal{D}_{n}$, thus completing the closed-loop training.